OpenStudy (anonymous):

How many no. of molecules of O2 gas are required to burn 10dm^3 of C2H4 (ethene gas) at S.T.P ? C2H4+3O2 --------> 2CO2 + 2H2O

OpenStudy (chmvijay):

can u tell molecular weight of C2H4 and O2 :)

OpenStudy (anonymous):

O2 = 32 & C2H4 = 28

OpenStudy (chmvijay):

22400 ml gas =1 mole 10 ml of gas =1*10 /22400 = X (find X) moles of C2H4 1 mole of C2H4 requires 3 moles of O2 X moles requires =3 * X =Y ( find Y) 1 mole of O2 =6.023*10^23 molecules of O2 Y mole of O2 = 6.023 * Y =Z ( find Z ) its ur answer